# A kernel conditional two-sample test

### v1

@@ 3 Conditional testing
The general goal of conditional testing is to decide based on data whether a specific statement about a conditional distribution holds at any user-specified conditioning location $x\in\mathcal{X}$, which we call the covariate (parameter). Formally, a (conditional) hypothesis is a map $H:\mathcal{X}\times\Theta\to\{0,1\}$ such that, for every $p\in\Theta$, the fulfillment region of $H$ for $p$, defined as $\varphi_{H}(p):=\{x\in\mathcal{X}\mid H(x,p)~{}\text{holds}\}$, is measurable. Hypothesis testing then involves two hypotheses $H_{0}$ and $H_{1}$, called the null and alternative hypotheses, respectively. We assume throughout that $H_{1}(x,p)=\neg H_{0}(x,p)$, where $\neg$ denotes the logical negation. A conditional test is then a measurable map $\mathcal{T}:\mathcal{X}\times\mathcal{D}\to\{0,1\}$, interpreted as $\mathcal{T}(x,D)=1$ if, and only if, we reject the null hypothesis $H_{0}$ at the covariate $x\in\mathcal{X}$ based on the data $D\in\mathcal{D}$.
@@
The covariate rejection region is the subset of $\mathcal{X}$ on which the null hypothesis is rejected based on the data. It is defined for all $D\in\mathcal{D}$ as
@@
We emphasize that this definition is finer-grained than most related approaches [2, 3, 4, 5, 6, 7, 8], which correspond to taking $\max_{\mathcal{X}}\mathcal{T}(\cdot,D)$ for the outcome of the test. A conditional test has a “perfect” outcome when the covariate rejection region $\chi(D)$ exactly recovers the complement of the fulfillment region of $H_{0}$, which satisfies $\varphi_{H_{0}}(p)^{C}=\varphi_{H_{1}}(p)$. It follows that a test can make two “types” of errors at a covariate $x\in\mathcal{X}$, corresponding respectively to whether the covariate belongs to the sets
@@
For $(\mathcal{S},\mathcal{N})\in\mathcal{A}_{\mathcal{X}}\times 2^{\mathbb{N}_{>0}}$, we say that the test $\mathcal{T}$ has type $i$ $(\mathcal{S},\mathcal{N})$-guarantee $\alpha\in(0,1)$ if $\sup_{p\in\Theta}\eta_{i}(p,\mathcal{S},\mathcal{N})\leq\alpha$. A type $1$ $(\mathcal{S},\mathcal{N})$-guarantee is called an $(\mathcal{S},\mathcal{N})$-level. Finally, if $\mathcal{S}=\{x\}$ or $\mathcal{N}=\{n\}$, we abuse notation in the above definitions and replace the subset by its only element.
@@
We emphasize the difference between having an $(x,\cdot)$-guarantee for all $x\in\mathcal{X}$ and an $(\mathcal{X},\cdot)$-guarantee. The former allows using the test at a single arbitrary covariate, whereas the latter guarantees the performance when using the test jointly on arbitrary covariates and enables trusting the rejection region as a whole. Similarly, $(\cdot,n)$-guarantees for all $n\in{\mathbb{N}_{>0}}$ allow for a one-time use of the test at an arbitrary time $n\in{\mathbb{N}_{>0}}$, whereas a $(\cdot,{\mathbb{N}_{>0}})$-guarantee allows for sequential testing. Our tests have $(\mathcal{X},{\mathbb{N}_{>0}})$- or $(\mathcal{X},n)$-guarantees, depending on assumptions.
Whether error functions are low at certain covariates depends on the null hypothesis, prior set, and family of data-generating processes. Consider the case where $\mathcal{X}=\{x_{1},x_{2}\}$, $X_{p,n}=x_{1}$ a.s. for all $p\in\Theta$, and $\Theta$ is such that $p(\cdot,x_{1})$ and $p(\cdot,x_{2})$ can be chosen independently. Since the null hypothesis should correspond to the “uninformed guess” — meaning that we only wish to reject it if we have significant evidence speaking against it — we should never reject the null hypothesis in $x_{2}$, due to the absence of data at that point. Thus, $\eta_{\mathrm{II}}(p,x_{2},\mathcal{N})=1$ for any Markov kernel such that $H_{0}$ does not hold in $x_{2}$. More generally, a low type $\mathrm{II}$ error function out of samples can only result from prior assumptions, encoded in $\Theta$. These concerns are void for the type $\mathrm{I}$ error function since the covariate rejection should only contain covariates where data suffices to reject the null hypothesis. This asymmetry of “reasonable expectations” between the type $\mathrm{I}$ and $\mathrm{II}$ errors thus results directly from properties of the test $\mathcal{T}$ rather than from the setup, which is symmetric between $H_{0}$ and $H_{1}$.
4 Conditional two-sample tests
@@
We specialize the preceding developments to conditional two-sample tests with finite-sample guarantees. We are interested in testing the null hypothesis [^3] Formally, $H_{0}(x,p_{1},p_{2})=1$ if, and only if, $\mathbb{E}(p_{1})(x)=\mathbb{E}(p_{2})(x)$.
$H_{0}(x,p_{1},p_{2}):\mathbb{E}(p_{1})(x)=\mathbb{E}(p_{2})(x),$ (5)
@@
where $(p_{1},p_{2})\in\Theta_{1}\times\Theta_{2}$ and $\Theta_{i}$ is a set of Markov kernels from $\mathcal{X}$ to $\mathcal{Y}$ with first moments, $i\in\{1,2\}$. First, we provide an abstract result to go from confidence bounds of a learning algorithm to a conditional two-sample test (Theorem 4.2). We then specialize it to KRR in the well-specified case, that is, assuming that the conditional expectations lie in the RKHS, effectively obtaining two kernel conditional two-sample tests which differ in their assumptions on the data generation and the time intervals on which guarantees hold (Theorem 4.4). The guarantees are based on new confidence bounds for KRR, which we present first and are of independent interest (Theorem 4.3). In practice, we leverage the bootstrapping scheme detailed in Appendix B to obtain sharper thresholds and sidestep the tuning of some inaccessible parameters. We also remind the reader interested in practical summaries and insights that Appendix A is a tutorial showcasing the tools developed in this section to compare the moments of conditional distributions.
@@
The conditional two-sample hypothesis of (5) can be equivalently reformulated into the setup of Section 3, but we forego doing it explicitly for brevity. We emphasize, however, that we abuse notation and redefine the test $\mathcal{T}$, covariate rejection region $\chi$, error regions $E_{i}$, and error functions $\eta_{i}$, $i\in\{\mathrm{I},\mathrm{II}\}$, to take as inputs (where applicable) two data sets $D_{1}$ and $D_{2}$, two Markov kernels $p_{1}$ and $p_{2}$, and two sets of integers of interest $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$. Specifically, they are now defined as $\chi(D_{1},D_{2})=\mathcal{T}(\cdot,D_{1},D_{2})^{-1}(\{1\})$, $E_{\mathrm{I}}(p_{1},p_{2},D_{1},D_{2})=\chi(D_{1},D_{2})\cap\varphi_{0}(p_{1},p_{2})$, $E_{\mathrm{II}}(p_{1},p_{2},D_{1},D_{2})=\chi(D_{1},D_{2})^{C}\cap\varphi_{1}(p_{1},p_{2})$, and
@@ 4.1 From confidence bounds to hypothesis testing
Let $\mathcal{T}:=\mathcal{T}_{(B_{1},B_{2})}$ be the conditional two-sample test defined as
@@
Then, $\mathcal{T}$ has $(\mathcal{S},\mathcal{N}_{1},\mathcal{N}_{2})$-level $\delta_{1}+\delta_{2}$ when used on the families $D^{(1)}$ and $D^{(2)}$.
The proof is in Appendix E. In more intuitive terms, this theorem guarantees that if one has two learning methods that come with associated confidence bounds (6) for a specific data-generating process (given by the functions $B_{i}$) uniformly in $\Theta$, then one can leverage these intervals for a conditional two-sample test with (covariate-dependent) statistic $\lVert f_{D_{1}}^{(1)}(x)-f_{D_{2}}^{(2)}(x)\rVert$. This result is very natural, and its value resides in clarifying the simple relationship between estimation and testing.
4.2 A kernel conditional two-sample test
@@ 4.2.1 Confidence bounds for vector-valued KRR
This result follows from Theorem D.23, which provides confidence bounds for vector-valued least-squares (not necessarily in \@iaciRKHS RKHS) under general assumptions. Theorem 4.3 and its generalization Theorem D.23 are — together with the general setup of Section 3 — the main theoretical contributions of this work. A detailed discussion of the connections between Theorems 4.3 and D.23 and similar bounds is available in Appendix C.
4.2.2 Application to two-sample testing
We can immediately apply the previous bounds to obtain kernel-based conditional two-sample tests.
@@ 4.2.2 Application to two-sample testing
    the assumptions of Case 1 in Theorem 4.3 hold for $R_{i}$ and $K$. Then, define $\mathcal{N}_{i}={\mathbb{N}_{>0}}$, and take $B_{i}(D,x)=\beta_{i}(D,\alpha_{i})\cdot\sigma_{D,\lambda_{i}}(x)$, for all $(D,x)\in\mathcal{D}\times\mathcal{X}$, where $\beta_{i}:=\beta_{\lambda_{i}}$ is defined in (10) with the corresponding choice of parameters;
@@
    the assumptions of Case 2 in Theorem 4.3 hold for $R_{i}$ and $D^{(i)}_{p}$ for all $p\in\Theta$. Then, define $\mathcal{N}_{i}=\{N_{i}\}$ for some $N_{i}\in{\mathbb{N}_{>0}}$, and take $B_{i}(D,x)=\beta_{i}(D,\alpha_{i})\cdot\sigma_{D,\lambda_{i}}(x)$, for all $(D,x)\in\mathcal{D}\times\mathcal{X}$, where $\beta_{i}:=\beta_{\lambda_{i}}$ is defined in (11) with the corresponding choice of parameters;
Then, with this choice for $B_{i}$ in (6) for all $i\in\{1,2\}$, the test (7) has $(\mathcal{X},\mathcal{N}_{1},\mathcal{N}_{2})$-level $\alpha_{1}+\alpha_{2}$ when used on the families $D^{(1)}$ and $D^{(2)}$.
This result formalizes the natural idea of comparing the norm difference between outcomes of KRR ran on both data set to the sum of the (appropriately scaled) posterior variances of the associated Gaussian processes. The test is not fully distribution-free [1], but allows data generated online. In practice, we leverage the bootstrapping scheme described in Appendix B to avoid tuning $\beta_{i}$.
Figure 1: Analytical and bootstrapped tests on a toy example. The tests reject $H_{0}$ (red thick line) when the confidence intervals (shaded regions) do not overlap, and accept it (green thick line) otherwise.
@@ 5 Numerical results
We begin with an example illustrating the different components. We then evaluate performance (type $\mathrm{I}$ and $\mathrm{II}$ errors) in controlled settings compared to the baseline of Hu and Lei [5]. Next, we illustrate benefits of our test and its pointwise answers compared to global conditional tests, thanks to the covariate rejection region and the fact that we obtain lower type $\mathrm{II}$ error when the tested functions differ only in rarely-sampled covariate regions. We remind the reader that Appendices A and B contain further numerical studies respectively showcasing how to compare moments of conditional distributions and perform monitoring, and validating our bootstrapping scheme. Finally, we emphasize that the purpose of this numerical study is not to establish dominance over other methods, but to investigate numerically how the framework works in well-understood scenarios. Additional information on experiments, hyperparameters, and the link to the code are in Appendix F.
@@
We refer below to “type $\mathrm{I}$ or $\mathrm{II}$ errors” and “positive rates” without referring to the corresponding triple $(\mathcal{S},\mathcal{N}_{1},\mathcal{N}_{2})$. They are meant as $(\mathcal{S},n,n)$-errors of type $\mathrm{I}$ or $\mathrm{II}$  and as the positive rate of the test triggering anywhere in the subset of interest $\mathcal{S}$ with data sets of lengths $n\in{\mathbb{N}_{>0}}$. We approximate triggering anywhere in $\mathcal{S}$ by evaluating the test at the locations present in the data set. A trigger is a type $\mathrm{II}$ error if it occurs in regions where $H_{0}$ is not enforced. We say below that we “pick” a function to indicate that we randomly choose it from the unit sphere of the RKHS $\mathcal{H}_{K}$ on $\mathcal{X}$. We take $K$ as the scalar Gaussian kernel with bandwidth $\gamma^{2}=0.25$ [26]; see Appendix A for non-scalar-valued $K$. Unless stated, data sets have $100$ points and are sampled uniformly.
@@
We now set $\mathcal{X}=[-3,3]^{2}$, pick $f_{1}$, and set $f_{2}(\cdot)=k(\cdot,0)$. We generate data from $f_{1}$ and $f_{1}+f_{2}$, which differ on $\mathcal{X}_{\mathrm{diff}}\coloneqq\{x\in\mathcal{X}\mid f_{2}(x)\geq 10^{-2}\}$ and (approximately) coincide on $\mathcal{X}_{\mathrm{same}}\coloneqq\mathcal{X}\setminus\mathcal{X}_{\mathrm{diff}}$. Data is sampled via a mixture of uniform distributions on $\mathcal{X}_{\mathrm{same}}$ and on $\mathcal{X}_{\mathrm{diff}}$. Figure 2 (right) shows the type $\mathrm{II}$ error as a function of $\theta$, the weight of $\mathcal{X}_{\mathrm{diff}}$ in the mixture. Densities are not estimated but set to their ground truth for the baseline [5]. Our test achieves higher power for low values of $\theta$. This directly results from the locality of our tests, whereas the baseline computes one aggregated statistic. This demonstrates the advantage of local guarantees.

### v2

@@ 3 Conditional testing
The general goal of conditional testing is to decide based on data whether a specific statement about a conditional distribution holds at any user-specified conditioning location $x\in\mathcal{X}$, which we call the covariate (parameter). Formally, a (conditional) hypothesis is a map $H:\mathcal{X}\times\Theta\to\{0,1\}$ such that, for every $p\in\Theta$, the fulfillment region of $H$ for $p$, defined as $\varphi_{H}(p):=\{x\in\mathcal{X}\mid H(x,p)=1\}$, is measurable. Hypothesis testing then involves two hypotheses $H_{0}$ and $H_{1}$, called the null and alternative hypothesis, respectively. We assume throughout that $H_{1}(x,p)=\neg H_{0}(x,p)$, where $\neg$ denotes the logical negation. A conditional test is then a measurable map $\mathcal{T}:\mathcal{X}\times\mathcal{D}\to\{0,1\}$, interpreted as $\mathcal{T}(x,D)=1$ if, and only if, we reject the null hypothesis $H_{0}$ at the covariate $x\in\mathcal{X}$ based on the data $D\in\mathcal{D}$.
@@
The covariate rejection region of a conditional test $\mathcal{T}$ is the subset of $\mathcal{X}$ on which the null hypothesis is rejected based on the data. It is defined for all $D\in\mathcal{D}$ as
@@
We emphasize that this definition is finer-grained than most related approaches [2, 3, 4, 5, 6, 7, 8], which correspond to taking $\max_{\mathcal{X}}\mathcal{T}(\cdot,D)$ for the outcome of the test. A conditional test has a “perfect” outcome when the covariate rejection region $\chi(D)$ exactly recovers the complement of the fulfillment region of $H_{0}$, which satisfies $\varphi_{0}(p)^{C}=\varphi_{1}(p)$ and where we introduced the shorthands $\varphi_{i}:=\varphi_{H_{i}}$, $i\in\{0,1\}$. It follows that a test can make two “types” of errors at a covariate $x\in\mathcal{X}$, corresponding respectively to whether the covariate belongs to the sets
@@
For $(\mathcal{S},\mathcal{N})\in\mathcal{A}_{\mathcal{X}}\times 2^{\mathbb{N}_{>0}}$, we say that the test $\mathcal{T}$ has type $i$ $(\mathcal{S},\mathcal{N})$-guarantee $\alpha\in(0,1)$ if $\sup_{p\in\Theta}\eta_{i}(p,\mathcal{S},\mathcal{N})\leq\alpha$. A type $\mathrm{I}$ $(\mathcal{S},\mathcal{N})$-guarantee is called an $(\mathcal{S},\mathcal{N})$-level. Finally, if $\mathcal{S}=\{x\}$ or $\mathcal{N}=\{n\}$, we abuse notation in the above definitions and replace the subset by its only element.
@@
We emphasize the difference between having an $(x,\cdot)$-guarantee for all $x\in\mathcal{X}$ and an $(\mathcal{X},\cdot)$-guarantee. The former allows using the test at a single arbitrary covariate, whereas the latter guarantees the performance when using the test jointly on arbitrary covariates and enables trusting the rejection region as a whole. Similarly, $(\cdot,n)$-guarantees for all $n\in{\mathbb{N}_{>0}}$ allow for a one-time use of the test at an arbitrary time $n\in{\mathbb{N}_{>0}}$, whereas a $(\cdot,{\mathbb{N}_{>0}})$-guarantee allows for sequential testing. Our tests have $(\mathcal{X},{\mathbb{N}_{>0}})$- or $(\mathcal{X},n)$-guarantees, depending on which assumptions are made.
While the formal definitions of type $\mathrm{I}$ and type $\mathrm{II}$ error functions are symmetric, their reasonable targets are not. Type $\mathrm{I}$ errors are those we seek to control; their frequency must be less than the prescribed level everywhere in $\mathcal{S}$, including in regions with little data. A false rejection of $H_{0}$ must not be caused by lack of information. As a result, control of type $\mathrm{II}$ errors is only achievable in regions of the covariate space where samples provide sufficient information (either through local sample mass or extrapolation enabled by the prior set). In other words, the type $\mathrm{II}$ error function can only be low where the data actually enable distinguishing the alternative from the null.
4 Testing of conditional expectations and functionals
@@
We specialize the preceding developments to tests of conditional expectations with finite-sample guarantees. We are interested in testing the null hypothesis [^3] Formally, $H_{0}(x,p_{1},p_{2})=1$ if, and only if, $\mathbb{E}(p_{1})(x)=\mathbb{E}(p_{2})(x)$.
$H_{0}(x,p_{1},p_{2}):\mathbb{E}(p_{1})(x)=\mathbb{E}(p_{2})(x)\quad\text{v.s.}\quad H_{1}(x,p_{1},p_{2}):\mathbb{E}(p_{1})(x)\neq\mathbb{E}(p_{2})(x),$ (5)
@@
where $(p_{1},p_{2})\in\Theta_{1}\times\Theta_{2}$ and $\Theta_{i}$ is a set of Markov kernels from $\mathcal{X}$ to $\mathcal{Y}$ with first moments, $i\in\{1,2\}$. First, we provide an abstract result to go from confidence bounds of a learning algorithm to a test of conditional expectations (Theorem 4.2). We then specialize it to KRR in the well-specified case, obtaining two tests (Theorem 4.4) with similar thresholds but differing guarantees and assumptions; they are based on new confidence bounds for KRR which we present first (Theorem 4.3). We finally show how to specialize the test to functionals other than the expectation, allowing (among others) a conditional two-sample test. This last point is further developed in Appendix A.
@@
The hypothesis (5) can be equivalently reformulated into the setup of Section 3, but we forego doing it explicitly for brevity. We emphasize, however, that we abuse notation and redefine the test $\mathcal{T}$, covariate rejection region $\chi$, error regions $E_{i}$, and error functions $\eta_{i}$, $i\in\{\mathrm{I},\mathrm{II}\}$, to take as inputs (where applicable) two data sets $D_{1}$ and $D_{2}$, Markov kernels $p_{1}$ and $p_{2}$, and sets of integers of interest $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$. Specifically, they are now defined as $\chi(D_{1},D_{2})=\mathcal{T}(\cdot,D_{1},D_{2})^{-1}(\{1\})$, $E_{\mathrm{I}}(p_{1},p_{2},D_{1},D_{2})=\chi(D_{1},D_{2})\cap\varphi_{0}(p_{1},p_{2})$, $E_{\mathrm{II}}(p_{1},p_{2},D_{1},D_{2})=\chi(D_{1},D_{2})^{C}\cap\varphi_{1}(p_{1},p_{2})$, and
@@ 4.1 From confidence bounds to hypothesis testing
Let $\mathcal{T}:=\mathcal{T}_{(B_{1},B_{2})}$ be the test defined as
@@
Then, $\mathcal{T}$ is a test of $H_{0}$ with $(\mathcal{S},\mathcal{N}_{1},\mathcal{N}_{2})$-level $\delta_{1}+\delta_{2}$ when used on the families $D^{(1)}$ and $D^{(2)}$.
The proof is in Appendix F. In more intuitive terms, this theorem guarantees that if one has two learning methods that come with associated confidence bounds (6) for a specific data-generating process (given by the functions $B_{i}$) uniformly in $\Theta$, then one can leverage these intervals for a conditional two-sample test with (covariate-dependent) statistic $\lVert f_{D_{1}}^{(1)}(x)-f_{D_{2}}^{(2)}(x)\rVert$. This result is very natural, and its value resides in clarifying the simple relationship between estimation and testing.
4.2 A kernel test of conditional expectations
@@ 4.2.1 Confidence bounds for vector-valued KRR
For the readers familiar with GP regression [36, 37, 26], this result guarantees that the graph of the ground truth lies in a tube centered at the GP posterior mean with width the posterior variance scaled by $\beta_{\lambda}$, with high probability. The bound is frequentist, however; the standard Bayesian assumption of a GP prior is replaced by the one that the conditional expectation lies in the RKHS. This result follows from Theorem E.23, which provides confidence bounds for vector-valued least-squares (not necessarily in \@iaciRKHS RKHS) under general assumptions. Theorem 4.3 and its generalization Theorem E.23 are — together with the general setup of Section 3 — the main theoretical contributions of this work. A detailed discussion of the connections between Theorems 4.3 and E.23 and similar bounds is available in Appendix D.
4.2.2 Application to testing
We can immediately apply the previous bounds to obtain kernel-based tests.
@@ 4.2.2 Application to testing
    The assumptions of Case 1 in Theorem 4.3 hold for $R_{i}$ and $K$. Then, define $\mathcal{N}_{i}={\mathbb{N}_{>0}}$, and take $B_{i}(D,x)=\beta_{i}(D,\alpha_{i})\cdot\sigma_{D,\lambda_{i}}(x)$, for all $(D,x)\in\mathcal{D}\times\mathcal{X}$, where $\beta_{i}:=\beta_{\lambda_{i}}$ is defined in (10) with the corresponding choice of parameters.
@@
    The assumptions of Case 2 in Theorem 4.3 hold for $R_{i}$ and $D^{(i)}_{p}$ for all $p\in\Theta$. Then, define $\mathcal{N}_{i}=\{N_{i}\}$ for some $N_{i}\in{\mathbb{N}_{>0}}$, and take $B_{i}(D,x)=\beta_{i}(D,\alpha_{i})\cdot\sigma_{D,\lambda_{i}}(x)$, for all $(D,x)\in\mathcal{D}\times\mathcal{X}$, where $\beta_{i}:=\beta_{\lambda_{i}}$ is defined in (11) with the corresponding choice of parameters.
Then, with this choice for $B_{i}$ in (6) for all $i\in\{1,2\}$, (7) is a test of $H_{0}$ with $(\mathcal{X},\mathcal{N}_{1},\mathcal{N}_{2})$-level $\alpha_{1}+\alpha_{2}$ when used on the families $D^{(1)}$ and $D^{(2)}$.
This result formalizes the natural idea of comparing the norm difference between outcomes of KRR ran on both data sets to the sum of the (appropriately scaled) posterior variances of the associated Gaussian processes. The test is not fully distribution-free [1] since it requires subgaussian Markov kernels, but allows data generated online. It also allows pointwise testing with guarantees — regardless of the input data distribution — thanks to the RKHS membership assumption. In practice, we leverage the bootstrapping schemes described in Appendix B to avoid tuning $\beta_{i}$.
Figure 1: Analytical and bootstrapped tests on a toy example. The tests reject $H_{0}$ (green thick line) when the confidence intervals (shaded regions) do not overlap, and accept it (red thick line) otherwise.
4.3 From conditional expectations to functionals
We now consider the announced case where we are interested in testing for equality of a functional $\mathcal{F}$ of the conditional distributions; that is, we are interested in a test of the null hypothesis
$H_{0}(x,p_{1},p_{2}):\mathcal{F}(p_{1}(\cdot,x))=\mathcal{F}(p_{2}(\cdot,x))\quad\text{v.s.}\quad H_{1}(x,p_{1},p_{2}):\mathcal{F}(p_{1}(\cdot,x))\neq\mathcal{F}(p_{2}(\cdot,x)).$ (12)
Examples involve $\mathcal{F}$ being moments of the distributions, or even the identity. Our core idea is to make (12) amenable to the abstract test of Theorem 4.2 by reformulating it as a test of conditional expectations. Specifically, we make the following assumption on $\mathcal{F}$. For consistency, we assume in this section that $p_{1}$ and $p_{2}$ are Markov kernels from $(\mathcal{X},\mathcal{A}_{\mathcal{X}})$ to a measurable set $(\mathcal{Z},\mathcal{A}_{\mathcal{Z}})$ which we call the measurement set, instead of mapping to $\mathcal{Y}$ directly.
Assumption 4.5.
There exists a separable Hilbert space $\mathcal{G}$ and a map $\Phi_{\mathcal{F}}:\mathcal{Z}\to\mathcal{G}$ called a representation map of $\mathcal{F}$ such that $\mathcal{Y}:=\Phi_{\mathcal{F}}(\mathcal{Z})\subset\mathcal{G}$ is closed, $\Phi_{\mathcal{F}}$ is Bochner-integrable w.r.t. any probability measure on $\mathcal{Z}$, and
$\forall(P,Q)\in{\mathcal{M}_{1}^{+}}(\mathcal{Z})^{2},\left(\mathcal{F}(P)=\mathcal{F}(Q)\quad\iff\quad\int_{\mathcal{Z}}\Phi_{\mathcal{F}}(z)\mathrm{d}P(z)=\int_{\mathcal{Z}}\Phi_{\mathcal{F}}(z)\mathrm{d}Q(z)\right).$
Under Assumption 4.5, (12) reduces to (5) where the Markov kernels involved are the pushforward kernels of $p_{1}$ and $p_{2}$ through $\Phi_{\mathcal{F}}$, which we denote with $q_{1}$ and $q_{2}$. Crucially, if a process of transition pairs $D^{(i)}:=((X_{n},Z_{n}))_{n\in{\mathbb{N}_{>0}}}$ of $p_{i}$ is available, the modified process $D_{\Phi_{\mathcal{F}}}^{(i)}:=((X_{n},\Phi_{\mathcal{F}}(Z_{n})))_{n\in{\mathbb{N}_{>0}}}$ is of transition pairs of $q_{i}$. In other words, any learning method with guarantees for $\mathcal{G}$-valued outputs can be used to test (12) via the data transformation $Y_{n}=\Phi_{\mathcal{F}}(Z_{n})$.
Assumption 4.5 is satisfied for many functionals of common interest, such as the ones extracting the moments of distributions or the identity — the latter makes (12) a conditional two-sample test. The test of Theorem 4.4 remains tractable even with infinite-dimensional $\mathcal{G}$: its complexity is exactly that of KRR. Furthermore, its assumptions such as subgaussianity are now on the pushforward kernels $q_{i}$ instead of the kernels $p_{i}$, $i\in\{1,2\}$. We detail this and provide numerical examples in Appendix A.
@@ 5 Numerical results
We begin with an example illustrating the different components. We then evaluate performance (type $\mathrm{I}$ and $\mathrm{II}$ errors) in controlled settings compared to the baseline of Hu and Lei [5]. Next, we illustrate benefits of our test and its pointwise answers compared to global conditional tests, thanks to the covariate rejection region and the fact that we obtain lower type $\mathrm{II}$ error when the tested functions differ only in rarely-sampled covariate regions. Finally, we showcase an application on change detection for a linear dynamical system. We remind the reader that Appendices A and B contain further numerical studies respectively comparing more general functionals (such as the two-sample one) and investigating our bootstrapping schemes. Finally, we emphasize that the purpose of this numerical study is not to establish dominance over other methods, but to investigate numerically how the framework works in well-understood scenarios. Additional information on experiments, hyperparameters, and the link to the code are in Appendix G.
@@
We refer below to “type $\mathrm{I}$ or $\mathrm{II}$ errors” and “positive rates” without referring to the corresponding triple $(\mathcal{S},\mathcal{N}_{1},\mathcal{N}_{2})$. They are meant as $(\mathcal{S},n,n)$-errors of type $\mathrm{I}$ or $\mathrm{II}$  and as the positive rate of the test triggering anywhere in the subset of interest $\mathcal{S}$ with data sets of lengths $n\in{\mathbb{N}_{>0}}$. We approximate triggering anywhere in $\mathcal{S}$ by evaluating the test at the locations present in the data set. A trigger is a type $\mathrm{II}$ error if it occurs in regions where $H_{0}$ is not enforced. We say we “pick” a function when we randomly choose it from the unit sphere of the RKHS $\mathcal{H}_{K}$ on $\mathcal{X}$. Except in the last experiment, $K$ is the scalar Gaussian kernel with bandwidth $\gamma^{2}=0.25$ [38]. Unless stated, data sets have $100$ uniformly-sampled points, and all results use the “naive” bootstrap (Algorithm 2).
@@
Figure 3: Left: Ratio between test statistic and threshold for the process monitoring example for various dimensions ($d$) and perturbation magnitudes ($\xi$). Triggers correspond to exceeding 11. Solid lines are means; shaded regions are $5\%$ and $95\%$ quantiles. Right: Ratio of $\sigma_{D,\lambda}(x)$ in (9) on the reference data $D$ between $x$ in the window and the initial state at $t=0$, averaged over the window.
@@
We now set $\mathcal{X}=[-3,3]^{2}$, pick $f_{1}$, and set $f_{2}(\cdot)=k(\cdot,0)$. We generate data from $f_{1}$ and $f_{1}+f_{2}$, which differ on $\mathcal{X}_{\mathrm{diff}}\coloneqq\{x\in\mathcal{X}\mid f_{2}(x)\geq 10^{-2}\}$ and (approximately) coincide on $\mathcal{X}_{\mathrm{same}}\coloneqq\mathcal{X}\setminus\mathcal{X}_{\mathrm{diff}}$. Data is sampled via a mixture of uniform distributions on $\mathcal{X}_{\mathrm{same}}$ and on $\mathcal{X}_{\mathrm{diff}}$. Figure 2 (right) shows the type $\mathrm{II}$ error as a function of $\theta$, the weight of $\mathcal{X}_{\mathrm{diff}}$ in the mixture. Densities are not estimated but set to their ground truth for the baseline [5]. Still, our test achieves higher power for low values of $\theta$. This directly results from the locality of our tests (whereas the baseline computes one aggregated statistic), which is enabled by the local guarantees of learning-theoretic bounds.
Process monitoring
We illustrate how the test allows dependent sampling by testing on the trajectories of a linear dynamical system on $\mathcal{X}=\mathbb{R}^{d}$; the goal is to detect a change in the dynamics. We consider a fixed reference data set of the transition pairs of $5$ trajectories of length $400$ with a fixed initial state. We compare it to a sliding window of length $50$ of the transition pairs of the current trajectory by performing the test at every point in the window. We perturb the dynamics after $200$ steps with a magnitude controlled monotonically by a scalar parameter $\xi$. The results are shown in Fig. 3. We successfully detect the change reliably, even before the window is completely filled with data from the perturbed dynamics. This supports that the test performs well despite correlations in sampling locations. We hypothesize based on Fig. 3 (right) and the interpretation of $\sigma_{D,\lambda}(x)$ as \@iaciGP GP posterior variance [37] that the degraded performance for a fixed $\xi$ as $d$ increases is mainly due to the perturbed dynamics driving the system in regions underexplored before the perturbation.